### v2

# Fragmentation of nuclear remnants in electron-nucleus collisions at high energy

Abstract: It is widely accepted that the various internal structures of an excited nucleus can lead to different topological configurations of nuclear fragments during fragmentation. However, these internal structures may transition from one configuration to another with varying probabilities. Utilizing a partitioning method based on equal (or unequal) probabilities—without incorporating the alpha-cluster ($\alpha$-cluster) model—allows for the derivation of diverse topological configurations of nuclear fragments resulting from fragmentation. Subsequently, we predict the multiplicity distribution of nuclear fragments for specific excited nuclei, such as 9Be, 12C, and 16O, which can be formed as nuclear remnants in electron-nucleus ($eA$) collisions at high energy. According to the $\alpha$-cluster model, an $\alpha$-cluster structure should take precedence among different internal structures; this may result in deviations in the multiplicity distributions of nuclear fragments with charge $Z=2$, compared to those predicted by the partitioning methods. The findings presented herein could serve as a reference for validating the existence of $\alpha$-cluster structures.
Keywords: $\alpha$-cluster structure; nuclear structure; multiplicity distribution of fragments; the Electron-Ion Collider
PACS Number(s): 21.60.Gx, 21.60.-n, 21.60.Ka
1 Introduction
High-energy nuclear collisions represent a significant area of research within modern physics [2, 3, 4, 5, 6]. In these interactions, numerous particles are predominantly generated within participant regions while multiple fragments are primarily emitted from spectator regions when available. During multi-fragment emission processes, it is anticipated that spectators will form an excited nucleus. This excited nucleus subsequently undergoes fragmentation into various components [7, 8, 9, 10]. Such fragmentation reveals rich internal structures within the nucleus where protons and neutrons can combine appropriately to create intermediate configurations [11, 12, 13, 14]. When two protons and two neutrons coalesce into such an intermediate structure, it is referred to as an alpha-cluster ($\alpha$-cluster) structure [15, 16, 17, 18]. It is possible that two or more $\alpha$-cluster structures are existent in a heavy nucleus [19, 20, 21, 22, 23]. According to the $\alpha$-cluster model [24, 25, 26, 27, 28], the $\alpha$-cluster structure should have much higher probability than other intermediate structures.
Different configurations of nuclear fragments can be measured in the fragmentation of excited nuclei. A multi-$\alpha$ configuration is one such arrangement that arises from statistical or stochastic fragmentation processes. Several experimental results concerning multi-$\alpha$ configurations in 16O fragmentation at high energies have been reported in the literature [7, 29, 30, 31, 32]. According to the $\alpha$-cluster model [24, 25, 26, 27, 28], further evidence and a higher probability for the presence of $\alpha$-cluster structures are anticipated. If the occurrence of multi-$\alpha$ configurations is significantly more probable than what would be expected based on partitioning methods derived from stochastic processes, it may indeed reflect the underlying $\alpha$-cluster structure of the excited nucleus.
Due to challenges in excluding the influence of partitioning probabilities across various configurations, there remains limited experimental evidence supporting $\alpha$-cluster structures; this evidence is insufficient for a comprehensive validation of the $\alpha$-cluster model [24, 25, 26, 27, 28]. Consequently, there is a pressing need for more systematic experimental investigations. This necessity motivates researchers to measure fragmentation products originating from excited nuclei. It is anticipated that diverse types of excited nuclei (nuclear remnants) will be formed through nuclear reactions induced by electrons at the forthcoming Electron-Ion Collider (EIC) [33]. The EIC presents an exceptional opportunity for researchers to systematically explore $\alpha$-cluster structures and thoroughly validate the $\alpha$-cluster model using fragmentation products from excited nuclei.
In this work, both equal and unequal probability partitioning methods are employed to derive various configurations of nuclear fragments from excited states such as 9Be, 12C, and 16O, which are expected to result from electron-nucleus ($eA$) collisions at the EIC [33]. The multiplicity distributions for all fragments as well as those with charge $Z$ will subsequently be obtained. In particular, the probability of multi-$\alpha$ or multi-He configuration or channel can be obtained, which may serve as the baseline for judging about the $\alpha$-cluster structure.
2 Various configurations of nuclear fragments
In the context of multi-fragment emission during $eA$ collisions, various fragmentation properties warrant special attention. For instance, understanding the types of fragments is crucial for elucidating the mechanisms underlying nuclear fragmentation; however, determining the number of neutrons in an isotope presents a complex challenge. Our previous research has demonstrated that the isotopic production cross section follows an Erlang distribution [34]. When different isotopes with a given charge number are not distinguished from one another, the analysis becomes significantly more straightforward.
At the EIC, as nuclear remnants, excited nuclei formed in $eA$ collisions can fragment into diverse topological configurations. This allows for an investigation into the internal structure of these excited nuclei. During fragmentation, both proton and neutron numbers are conserved. In experimental settings, it is possible to measure either the charge or proton count of a fragment. This capability facilitates our examination of multiplicity distributions among fragments with varying charges. Furthermore, we may delve deeper into discussing the fundamental physical reasons behind these multiplicity distributions observed in nuclear fragments. Notably, factors such as $\alpha$-cluster structures and liquid-gas phase transitions could influence these experimentally measured multiplicity distributions.
As examples, we now consider three types of $eA$ collisions at the EIC:
[…]
and
[…]
Then, the excited 9Be, 12C, and 16O nuclei can be obtained and analyzed. Other excited nuclei are not the focus of the present work due to the fact that they do not have an advantage in the study of $\alpha$-cluster structure.
It is important to note that the process $e+n$ or $e+p$ occurring in $eA$ is not electron-induced neutron/proton knock-out reaction, which typically manifests at beam energies of hundreds of MeV [35, 36]. Instead, this represents multi-particle production process that occurs at beam energies on the order of hundreds of GeV, for which the EIC is specifically designed to achieve [37], with a center-of-mass energy range between 20 and 100 GeV. In the incident nucleus $A$, alongside the participant nucleon, there exist spectator nucleons—the remaining constituents—which will form an excited nucleus characterized by energy levels significantly higher than those attainable through MeV collisions.
The excited nuclei subsequently decay into various nuclear fragments. The correlations between momentum and scattering angle for evaporated neutrons and protons have been extensively studied using the BeAGLE (Benchmark eA Generator for LEptoproduction) model [38], particularly in high-energy lepton-nucleus collisions. The excited nuclei produced in $eA$ collisions at the EIC exhibit significantly higher excitation levels compared to those generated in electron-induced neutron/proton knock-out reactions conducted with fixed targets at low and medium energies. Due to substantial excitation leading to large internal momenta, both decay protons and other nuclear fragments correspond to sufficiently large polar angles that fall within the estimated pseudorapidity acceptance region designated for the currently proposed EIC detector [33, 39].
Considering the Fermi momentum of a nucleon within the nucleus, which is approximately 0.25 GeV$/c$, and the momentum per nucleon of the incident nucleus being 10 GeV$/c$, decay protons and other fragments are expected to be emitted within a forward cone characterized by a polar angle $\theta_{0}=25$ mrad. This corresponds to a pseudorapidity of $\eta=-\ln\tan(\theta_{0}/2)=4.38$. Furthermore, recoil protons can be distinguished from decay protons since recoil protons participate in multi-scattering processes which result in much larger scattering angles. It is assumed that this emission will span a wide range from nearly 0 (corresponding to $\eta=\infty$) up to approximately $10\theta_{0}$ (which corresponds to $\eta=2.07$). Indeed, it cannot be excluded that some recoil protons may have very small scattering angles, leading to exceptionally large pseudorapidities due to the influence of leading nucleons.
If the two types of protons are assumed to emit isotropically in their respective rest frames, they approximately follow Gaussian $\eta$ distributions with a common standard deviation ($\sigma_{\eta}\approx 0.91$). The decay protons predominantly distribute within the range of $8.024.38<\eta<4.38+4\sigma_{\eta}=$ 8.02, while the recoil protons primarily occupy the range of $5.712.07<\eta<2.07+4\sigma_{\eta}=$ 5.71. It is evident that there exists some overlap between decay and recoil protons in the forward cone. Although most recoil protons may have emission angles exceeding 25 mrad due to multiple scattering processes, we cannot entirely dismiss the possibility of them appearing within the forward cone. To more effectively distinguish between decay and recoil protons, one could study their energies; generally, the energy of a decay proton is nearly equal to that of an incident nucleus per nucleon, whereas the energy of a recoil proton should be lower.
While it is challenging to precisely separate decay from recoil protons in the forward cone, such distinction is not essential for this study. In fact, among the three types of $eA$ collisions considered in previously mentioned reactions (1)–(3), our selected samples should ideally consist solely of those with only recoil neutrons; thus, any contributions from recoil protons must be excluded from our analysis. In rare instances where mixed events occur involving recoiling protons within the forward cone, these can introduce minor measurement errors. In cases where recoiling protons do appear in this region, misidentified events would include both these recoils and fragmentations from ${}^{9}{\rm Li}^{*}$, ${}^{12}{\rm B}^{*}$, and ${}^{16}{\rm N}^{*}$ respectively—these should be excluded from expected events associated with fragmentations originating from ${}^{9}{\rm Be}^{*}$, ${}^{12}{\rm C}^{*}$, and ${}^{16}{\rm O}^{*}$. Although there are clear differences in energy between decay and recoil protons, attempting to separate them experimentally may incur additional costs.
In our studies concerning nuclear fragments, our primary focus lies on counting electric charges rather than mass, momentum, energy, etc. The resolutions of detectors regarding secondary quantities do not impact our analysis; however, it is crucial that detector resolution for charge number remains high—approximately $\sim 2\%$, which is generally achievable. During experiments, it is essential to select relevant decay events where the total charge number of various nuclear fragments precisely matches that of the incident nucleus $A$. In some events, due to distribution fluctuations, individual nuclear fragment has probability to emit with a very small polar angle, which is mixed with the beam and cannot be captured by the detector [33, 37, 39]. Naturally, these events should be removed from the analysis.
In addition to $N_{F}$ representing the multiplicity of all nuclear fragments, let $N_{Z}$ be the multiplicity of the fragments with charge $Z$. In an equal probability partitioning method, in which the $\alpha$-cluster model does not enter, the frequency of configuration $\{N_{Z}(Z)\}$, or the weight of partition $\{N_{Z}(Z)\}$, is considered to be the same which results in the same probability $f_{1}$. Various topological configurations of nuclear fragments in excited nuclear fragmentation can be obtained by the treatment of exhaustive enumeration.
In an unequal probability partitioning method [40, 41], in which the $\alpha$-cluster model does not enter either, the frequency of configuration $\{N_{Z}(Z)\}$, or the weight of partition $\{N_{Z}(Z)\}$, is considered to be the number of exchange
$M_{2}=\frac{Q!}{\prod\limits_{Z}N_{Z}(Z)!Z^{N_{Z}}},$ (4)
where $Q$ is the charge number of the excited nucleus, $Q!$ and $N_{Z}(Z)!$ represent factorial operations, and $M_{2}$ is the Cauchy number in the number theory. The normalization of $M_{2}$ is
$\sum\limits_{\{N_{Z}(Z)\}}M_{2}=Q!.$ (5)
The probability of configuration $\{N_{Z}(Z)\}$ is
$f_{2}=\frac{M_{2}}{\sum\limits_{\{N_{Z}(Z)\}}M_{2}}=\frac{1}{\prod\limits_{Z}N_{Z}(Z)!Z^{N_{Z}}}.$ (6)
The equal and unequal probability partitioning methods present distinct perspectives in the realm of physics. The equal probability partitioning method is grounded in the principle of equal probability, a fundamental assumption in statistical physics. This principle asserts that when a system is at equilibrium, provided there are no additional constraints beyond energy, volume, and particle number, the likelihood of the system occupying each microscopic state remains uniform. Conversely, the unequal probability partitioning method relies on the principle of unequal probability; this acknowledges that within a sampling survey, the chance of selecting any individual from a population may vary due to the interchangeability of identical particles.
Prior to implementing the partitioning methods, it is essential to highlight other applications that demonstrate their validity and rationality. In previous studies [42, 43, 44, 45, 46], these methods were employed to investigate excited nuclear fragmentation during nucleus-nucleus collisions at intermediate and high energies. Conditional moments and their normalized forms across various orders were introduced [42, 43] for examining critical behavior [47, 48]. It was observed that correlations and distributions derived from conditional moments of nuclear fragments [44, 45, 46] obtained through the partitioning technique [40, 41] align well with experimental data concerning excited nuclear fragmentation resulting from diffractive excitation (nuclear reaction) as well as electromagnetic dissociation [49, 50].
Using the equal (unequal) probability partitioning method, various topological configurations of nuclear fragments in excited 9Be, 12C, and 16O fragmentation are listed in Tables 1, 2, and 3, respectively, in which each configuration has an equal (unequal) probability $f_{1}$ ($f_{2}$). The multiplicity, $N_{F}$, of all fragments and the multiplicity, $N_{Z}$, of the fragments with given charge $Z$ in a defined configuration are shown separately.
In the equal probability partitioning method, the numbers of configurations, or fragmentation channels, in fragmentation of excited 9Be, 12C, and 16O nuclei are 5, 11, and 22, respectively. The fragment Be is artificially assumed by default with 50% probability to be the most unstable 8Be and in 50% of the cases to be (relative) stable isotope of Be. 8Be can decay into 2He, which is listed in brackets with fractions in the tables. In the unequal probability partitioning method, the numbers of exchanges in excited 9Be, 12C, and 16O fragmentation are 24, 720, and 40320, respectively. The fragment Be is assumed by default with a given chance ($\{M_{2}(2{\rm He})/[M_{2}(2{\rm He})+M_{2}({\rm Be})]=1/3\}$) to be 8Be which is unstable and can decay into 2He with given fractions.
Table 1: The multiplicity, $N_{F}$, of all fragments and the multiplicity, $N_{Z}$, of the fragments with charge $Z$ in various configurations in excited 9Be fragmentation, where only the charge conservation is considered in the fragmentation. In the equal probability partitioning method, the fragment Be is defaulted with 50% probability to be 8Be, and in the unequal probability partitioning method, the fragment Be is defaulted with a given chance $\{M_{2}(2{\rm He})/[M_{2}(2{\rm He})+M_{2}({\rm Be})]=1/3\}$ to be 8Be, where 8Be is unstable and can decay into 2He, which is listed in the bracket, and causes $N_{F}$ to $N_{F}+1$ and $N_{Z=2}$ to $N_{Z=2}+2$. Here, the changeable $N_{F}$ and $N_{Z=2}$ are shown in the table by $+1$ and $+2$, respectively. The probabilities $f_{1}$ and $f_{2}$ of each channel obtained by the equal and unequal partitioning methods are listed respectively.
$N_{F}$ $N_{Z=1}$ $N_{Z=2}$ $N_{Z=3}$ $N_{Z=4}$ Configuration $f_{1}$ (1/5) $f_{2}$ (1/24)
4 4 4H 1 1
3 2 1 2H+He 1 6
2 1 1 H+Li 1 8
2 2 2He 1 3
1 1 [Be 0.5 4
1+1 +2 (2He)] 0.5 2
Table 2: The multiplicity $N_{F}$ of all fragments and the multiplicity $N_{Z}$ of the fragments with charge $Z$ in various configurations in excited 12C fragmentation, where only the charge conservation is considered in the fragmentation. In the equal probability partitioning method, the fragment Be is defaulted with 50% probability to be 8Be, and in the unequal probability partitioning method, the fragment Be is defaulted with a given chance (1/3) to be 8Be, where 8Be is unstable and can decay into 2He, which is listed in the bracket, and causes $N_{F}$ to $N_{F}+1$ and $N_{Z=2}$ to $N_{Z=2}+2$. The probabilities $f_{1}$ and $f_{2}$ of each channel obtained by the equal and unequal partitioning methods are listed respectively.
$N_{F}$ $N_{Z=1}$ $N_{Z=2}$ $N_{Z=3}$ $N_{Z=4}$ $N_{Z=5}$ $N_{Z=6}$ Configuration $f_{1}$ (1/11) $f_{2}$ (1/720)
6 6 6H 1 1
5 4 1 4H+He 1 15
4 3 1 3H+Li 1 40
4 2 2 2H+2He 1 45
3 2 1 [2H+Be 0.5 60
3+1 2 +2 2H+(2He)] 0.5 30
3 1 1 1 H+He+Li 1 120
3 3 3He 1 15
2 1 1 H+B 1 144
2 1 1 [He+Be 0.5 60
2+1 1+2 He+(2He)] 0.5 30
2 2 2Li 1 40
1 1 C 1 120
Table 3: The multiplicity $N_{F}$ of all fragments and the multiplicity $N_{Z}$ of the fragments with charge $Z$ in various configurations in excited 16O fragmentation, where only the charge conservation is considered in the fragmentation. In the equal probability partitioning method, the fragment Be is defaulted with 50% probability to be 8Be, and in the unequal probability partitioning method, the fragment Be is defaulted with a given chance (1/3) to be 8Be, where 8Be is unstable and can decay into 2He, which is listed in the bracket, and causes $N_{F}$ to $N_{F}+1$ and $N_{Z=2}$ to $N_{Z=2}+2$. The probabilities $f_{1}$ and $f_{2}$ of each channel obtained by the equal and unequal partitioning methods are listed respectively.
$N_{F}$ $N_{Z=1}$ $N_{Z=2}$ $N_{Z=3}$ $N_{Z=4}$ $N_{Z=5}$ $N_{Z=6}$ $N_{Z=7}$ $N_{Z=8}$ Configuration $f_{1}$ (1/22) $f_{2}$ (1/40320)
8 8 8H 1 1
7 6 1 6H+He 1 28
6 5 1 5H+Li 1 112
6 4 2 4H+2He 1 210
5 4 1 [4H+Be 0.5 280
5+1 4 +2 4H+(2He)] 0.5 140
5 3 1 1 3H+He+Li 1 1120
5 2 3 2H+3He 1 420
4 3 1 3H+B 1 1344
4 2 1 1 [2H+He+Be 0.5 1680
4+1 2 1+2 2H+He+(2He)] 0.5 840
4 2 2 2H+2Li 1 1120
4 1 2 1 H+2He+Li 1 1680
4 4 4He 1 105
3 2 1 2H+C 1 3360
3 1 1 1 H+He+B 1 4032
3 1 1 1 [H+Li+Be 0.5 2240
3+1 1 +2 1 H+(2He)+Li] 0.5 1120
3 2 1 [2He+Be 0.5 840
3 2+2 2He+(2He)] 0.5 420
3 1 2 He+2Li 1 1120
2 1 1 H+N 1 5760
2 1 1 He+C 1 3360
2 1 1 Li+B 1 2688
2 2 [2Be 0.25 504
2+1 +2 1 (2He)+Be 0.5 504
2+1+1 +2+2 (2He)+(2He)] 0.25 252
1 1 O 1 5040
Table 4: The multiplicity distribution, $dn/dN_{F}$, of all fragments and the multiplicity distribution, $dn/dN_{Z}$, of the fragments with charge $Z$ in excited 9Be fragmentation in the equal (unequal) probability partitioning method, where the normalization is 5 (24), which is the number of total configurations (exchanges). In the table, $N_{x}$ denotes $N_{F}$ or $N_{Z}$.
$dn/dN_{x}$ $N_{x}=0$ $N_{x}=1$ $N_{x}=2$ $N_{x}=3$ $N_{x}=4$
$dn/dN_{F}$ 0 (0) 0.5 (4) 2.5 (13) 1 (6) 1 (1)
$dn/dN_{Z=1}$ 2 (9) 1 (8) 1 (6) 0 (0) 1 (1)
$dn/dN_{Z=2}$ 2.5 (13) 1 (6) 1.5 (5) 0 (0) 0 (0)
$dn/dN_{Z=3}$ 4 (16) 1 (8) 0 (0) 0 (0) 0 (0)
$dn/dN_{Z=4}$ 4.5 (20) 0.5 (4) 0 (0) 0 (0) 0 (0)
Table 5: The multiplicity distribution $dn/dN_{F}$ of all fragments and the multiplicity distribution $dn/dN_{Z}$ of the fragments with charge $Z$ in excited 12C fragmentation in the equal (unequal) probability partitioning method, where the normalization is 11 (720), which is the number of total configurations (exchanges).
$dn/dN_{x}$ $N_{x}=0$ $N_{x}=1$ $N_{x}=2$ $N_{x}=3$ $N_{x}=4$ $N_{x}=5$ $N_{x}=6$
$dn/dN_{F}$ 0 (0) 1 (120) 2.5 (244) 3 (225) 2.5 (115) 1 (15) 1 (1)
$dn/dN_{Z=1}$ 4 (265) 2 (264) 2 (135) 1 (40) 1 (15) 0 (0) 1 (1)
$dn/dN_{Z=2}$ 5.5 (405) 2.5 (195) 1.5 (75) 1.5 (45) 0 (0) 0 (0) 0 (0)
$dn/dN_{Z=3}$ 8 (520) 2 (160) 1 (40) 0 (0) 0 (0) 0 (0) 0 (0)
$dn/dN_{Z=4}$ 10 (600) 1 (120) 0 (0) 0 (0) 0 (0) 0 (0) 0 (0)
$dn/dN_{Z=5}$ 10 (576) 1 (144) 0 (0) 0 (0) 0 (0) 0 (0) 0 (0)
$dn/dN_{Z=6}$ 10 (600) 1 (120) 0 (0) 0 (0) 0 (0) 0 (0) 0 (0)
Table 6: The multiplicity distribution $dn/dN_{F}$ of all fragments and the multiplicity distribution $dn/dN_{Z}$ of the fragments with charge $Z$ in excited 16O fragmentation in the equal (unequal) probability partitioning method, where the normalization is 22 (40320), which is the number of configurations (exchanges).
$dn/dN_{x}$ $N_{x}=0$ $N_{x}=1$ $N_{x}=2$ $N_{x}=3$ $N_{x}=4$ $N_{x}=5$ $N_{x}=6$ $N_{x}=7$ $N_{x}=8$
$dn/dN_{F}$ 0 (0) 1 (5040) 3.25 (12312) 5 (12516) 5.25 (7301) 3 (2660) 2.5 (462) 1 (28) 1 (1)
$dn/dN_{Z=1}$ 7 (14833) 4 (14832) 4 (7420) 2 (2464) 2 (630) 1 (112) 1 (28) 0 (0) 1 (1)
$dn/dN_{Z=2}$ 9.25 (22449) 5.5 (11340) 4 (4494) 1.5 (1260) 1.75 (777) 0 (0) 0 (0) 0 (0) 0 (0)
$dn/dN_{Z=3}$ 15 (29120) 5 (8960) 2 (2240) 0 (0) 0 (0) 0 (0) 0 (0) 0 (0) 0 (0)
$dn/dN_{Z=4}$ 19.25 (34272) 2.5 (5544) 0.25 (504) 0 (0) 0 (0) 0 (0) 0 (0) 0 (0) 0 (0)
$dn/dN_{Z=5}$ 19 (32256) 3 (8064) 0 (0) 0 (0) 0 (0) 0 (0) 0 (0) 0 (0) 0 (0)
$dn/dN_{Z=6}$ 20 (33600) 2 (6720) 0 (0) 0 (0) 0 (0) 0 (0) 0 (0) 0 (0) 0 (0)
$dn/dN_{Z=7}$ 21 (34560) 1 (5760) 0 (0) 0 (0) 0 (0) 0 (0) 0 (0) 0 (0) 0 (0)
$dn/dN_{Z=8}$ 21 (35280) 1 (5040) 0 (0) 0 (0) 0 (0) 0 (0) 0 (0) 0 (0) 0 (0)
Figure 1: Multiplicity distributions of nuclear fragments with different charges in 9Be fragmentation. The solid (dashed) histograms represent the results from the equal (unequal) probability partitioning method. Panel (a) is for all fragments. Panels (b)–(e) are for the fragments with charge $Z=1$, 2, 3, and 4, respectively.
Figure 2: Multiplicity distributions of nuclear fragments with different charges in 12C fragmentation. The solid (dashed) histograms represent the results from the equal (unequal) probability partitioning method. Panel (a) is for all fragments. Panels (b)–(g) are for the fragments with charge $Z=1$, 2, …, and 6, respectively.
Figure 3: Multiplicity distributions of nuclear fragments with different charges in 16O fragmentation. The solid (dashed) histograms represent the results from the equal (unequal) probability partitioning method. Panel (a) is for all fragments. Panels (b)–(i) are for the fragments with charge $Z=1$, 2, …, and 8, respectively.
3 Multiplicity distributions of nuclear fragments
The $N_{F}$ distribution, $dn/dN_{F}$, and the $N_{Z}$ distribution, $dn/dN_{Z}$, in fragmentation of excited 9Be, 12C, and 16O nuclei are given in Tables 4, 5, and 6, respectively, where $n$ denotes the frequency of $N_{F}$ occurring. The multiplicity distributions $dn/dN_{F}$ and $dn/dN_{Z}$ are also the yield distributions of nuclear fragments. The normalization constants of $dn/dN_{F}$ and $dn/dN_{Z}$ in the partitioning methods are the numbers of configurations (exchanges).
The normalized multiplicity distributions, $(1/n)(dn/dN_{F})$ or $(1/n)(dn/dN_{Z})$, in excited 9Be, 12C, and 16O fragmentation are displayed in Figures 1, 2, and 3, respectively. The solid (dashed) histograms represent the results from the equal (unequal) probability partitioning method. Figures 1(a), 2(a), and 3(a) are for the multiplicity distributions of all fragments. The multiplicity distributions of the fragments with different $Z$ are shown in different panels, where $Z=1$–4 in Figures 1(b)–1(e), $Z=1$–6 in Figures 2(b)–2(g), and $Z=1$–8 in Figures 3(b)–3(i) are for excited 9Be, 12C, and 16O fragmentation, respectively. It is worth noting that the predicted frequency distributions in Figures 1–3 are precise numerical values (fractions) in each bin under a given scenario.
One can see from Tables 4–6 and Figures 1–3 that the multiplicity distributions of the fragments with given $Z$ from both the equal and unequal probability partitioning methods have a quick decreasing trend in most cases. The larger the $Z$, the closer the trends of the two results are. In the equal probability partitioning method, the priority of 2He channel in 9Be fragmentation is significant, and the priority of 3He (4He) channel in 12C (16O) fragmentation is not significant. In the unequal probability partitioning method, the three cases do not show an obvious priority. Figures 1(a), 2(a), and 3(a) demonstrate peaks around the intermediate multiplicity, which are naturally different from the multiplicity distribution of the fragments with given $Z$.
In addition, the multiplicity distribution of the fragments with $Z=2$ can be seen clearly. In particular, in the equal probability partitioning method, $(dn/dN_{Z=2})/5=1.5/5=30\%$ for 2He channel in excited 9Be fragmentation, $(dn/dN_{Z=2})/11=1.5/11=13.64\%$ for 3He channel in excited 12C fragmentation, and $(dn/dN_{Z=2})/22=1.75/22=7.95\%$ for 4He channel in excited 16O fragmentation. In the unequal probability partitioning method, the three values are $5/24\approx 20.8\%$, $45/720=6.25\%$, and $777/40320\approx 1.93\%$.
One can see that the difference between the two percentages from the equal and unequal probability partitioning methods in 9Be fragmentation is not too large, and that in 16O fragmentation is quite large. If the $\alpha$-cluster structure does exist in excited nuclei formed in $eA$ collisions at the EIC, one should observe much more multi-He configuration than these percentages (probabilities), which can be obtained from Tables 4–6 (Figures 1–3). Although there are some experimental reports on the $\alpha$-cluster structure of excited nucleus [7, 29, 30, 31, 32], the related percentage or fraction of multi-He configurations is significantly smaller than that obtained through partitioning methods due to events with multi-particle production included in the data sample [29, 30, 31, 32]. Thus, this fraction cannot be directly compared with partitioning results.
An experimental study using nuclear emulsion [51] found that the H+2He channel fraction in 10B fragmentation at a beam energy of $E_{\rm beam}=1$ GeV/nucleon is 78%. Based on this finding, we estimate that the 2He channel fraction in 9Be fragmentation at $E_{\rm beam}=1$ GeV/nucleon is approximately 78%, possibly slightly higher due to fewer fragmentation channels for 9Be compared to 10B. The inferred 2He channel fraction in excited 9Be fragmentation is estimated to be 2.6–3.8 times that from partitioning methods. As a non-conservative estimation, we set our judgment line for $\alpha$ clustering cases at twice the baseline percentages (probabilities) without $\alpha$ clustering.
In this context, we assume that the experimental percentage of multi-He events follows a Gaussian distribution with standard deviation $\sigma$, predominantly concentrated within the range [$0,4\sigma$] and centered around an expected value of $2\sigma$, which serves as our baseline. If the experimental percentage exceeds $4\sigma$, defined as twice the baseline and serving as our threshold for judgment, one can draw conclusions regarding the existence of 3He or $\alpha$ clustering with over 95% confidence level. According to this line, we conclude that 3He or $\alpha$ clustering exists in excited 9Be formed during peripheral collisions between 10B and nuclear emulsion at $E_{\rm beam}=1$ GeV/nucleon.
Experimental data on the fragmentation of 9C, 10C, and 11C at an energy of $E_{\rm beam}=1.2$ GeV/nucleon within nuclear emulsion show fractions for the 3He channel as follows: 15.2%, 5.3%, and 17.5%, respectively [52, 53, 54]. Additionally, experiments on 16O fragmentation at $E_{\rm beam}=3.65$ and 200 GeV/nucleon within nuclear emulsion indicate fractions for the 4He channel as 12.5% and 2.3% respectively [55]. The fraction of multi-He channels in the fragmentation of excited 9,11C (16O) formed at $E_{\rm beam}=1.2$ (3.65) GeV/nucleon is more than double that predicted by unequal probability partitioning, suggesting the presence of 3He or $\alpha$ clustering in these excited nuclei. However, for excited 10C formed at $E_{\rm beam}=1.2$ GeV/nucleon and excited 16O formed at 200 GeV/nucleon, the multi-He channel fractions do not exceed twice those from unequal probability partitioning, indicating a stochastic result rather than 3He or $\alpha$ clustering.
It should be noted that the errors in experimental data quoted here are not available in refs. [51, 52, 53, 54, 55]. According to the errors in data for other channels [51], the relative errors for the quoted data are estimated by us to be 15–21%. Generally, at $E_{\rm beam}=1.2$ GeV/nucleon, the fraction of the 3He channel in excited 10C fragmentation is significantly lower than that in excited 9,11C due to its even-even nature, which enhances its stability and reduces 3He or $\alpha$ clustering probabilities while increasing other fragmentation channels’ likelihoods. Additionally, original $\alpha$ clustering presented in excited 16O at 3.65 GeV/nucleon is disrupted by violent collisions at 200 GeV/nucleon. These collisions lead to multi-particle production and participant nucleons separating from 16O, making conditions for forming 4He clusters less favorable. Furthermore, higher excitation energy achieved with increased beam energy likely surpasses threshold energies needed for forming such clusters; thus higher-excited 16O fragments into multiple nucleons instead of favoring a 4He channel.
Based on the judgment line, the fractions of 2He (3He or 4He) channel in excited 9Be (12C or 16O) fragmentation should be higher than 60% (27.28% or 15.9%) if the equal probability partitioning method is considered, or, 41.6% (12.5% or 3.86%) if the unequal probability partitioning method is considered. Here, these percentages are obtained from twice the values shown in Figures 1(c), 2(c), and 3(c), respectively, according to the assumption of twice the baselines. One may note that excited 9Be shows a significant 2He frequency and excited 12C (16O) does not show obvious enhancement of 3He (4He). The reason is that 9Be has very few fragmentation channels totally, and 12C (16O) has relatively more fragmentation channels totally. The tables and figures presented in the present work can be regarded as a benchmark reference result in which the $\alpha$-cluster model does not enter. We look forward to the results of excited nuclear fragmentation at the forthcoming EIC experiments to study the fraction of multi-He configuration.
In addition, in the excited nucleus formed in $eA$ collisions, a liquid-gas phase transition may also occur. In the above discussions on nuclear fragmentation, the liquid-gas phase transition is not taken into account in the calculations. If the liquid-gas phase transition occurs, more light fragments should be produced, causing the distribution of light fragment multiplicity to deviate from the histogram in Figures 1–3, reducing the probability of low multiplicity events and increasing the probability of high multiplicity events. Meanwhile, heavy fragments should not be produced, or their yield should be very low. The results of this work can also provide reference for whether liquid-gas phase transition occurs in the excited nucleus in $eA$ collisions at the future EIC.
In $eA$ collisions, if the incident nucleus $A$ is very large, the liquid-gas phase transition can occur in a part of the excited nucleus. For the local area, where the phase transition has occurred, many light fragments are expected to be emitted, and there is no intermediate and heavy fragment emitted with them. For the remainder area, where the phase transition has not happened, the fragmentation is not special, in which the multiplicity distribution of nuclear fragments should generally obey the partitioning methods.
To ensure an accurate description as possible, the heaviest fragments produced in an event—considered remnants of excited nuclear fragmentation—should be excluded from analysis. For genuine evaporation products, it must be acknowledged that they originate from the fragmentation process involving a smaller excited nucleus; thus, the partitioning methods should be reapplied specifically for this smaller nucleus. Whether a liquid-gas phase transition occurs in the overall or local area, the proportion of light fragments with $Z=1$ should exceed twice the baseline values when applying Poisson distribution to the considered probabilities. Furthermore, if experimental measurements fall within the theoretical uncertainty range, fragmentation may be interpreted as a consequence of a general stochastic process.
Beyond $\alpha$ clustering and liquid-gas phase transitions—which may lead to significant deviations between experimental multiplicity distributions and theoretical models—other nuclear effects exert only minor influences on experimental outcomes. These nuclear effects involved include non-uniform nucleon number density distributions (the neutron skin structure of heavy nuclei), symmetrical energy characteristics of nuclear matter, two- or multi-nucleon correlations within nuclei, as well as stopping power or transparency phenomena associated with nuclear interactions. Here by the slight effects it is meant that both the effects themselves and their impact can be neglected in studying the multiplicity distribution of nuclear fragments.
The reason why other nuclear effects are small is that they mainly affect the momentum distribution of nucleons inside the nucleus. Due to limited strength, the other nuclear effects mentioned above are not sufficient to affect the formation of nuclear fragments with given charge $Z$, though they affect the neutron numbers in emitted isotopes. As a result, they also affect the kinetic energy and emitting direction of nuclear fragments. In short, the transverse momentums and polar angles of nuclear fragments are significantly affected, while the charges and multiplicity of nuclear fragments are slightly affected.
4 Summary
In summary, various configurations of nuclear fragments resulting from the fragmentation of excited 9Be, 12C, and 16O nuclei—expected to form in $eA$ collisions at the EIC—are investigated using both equal and unequal probability partitioning methods. The multiplicity distributions for all fragments as well as those with charge $Z$ are derived. In comparison to results obtained from these partitioning methods, experiments suggest that multi-$\alpha$ configurations should exhibit a significantly high probability according to the $\alpha$-cluster model. We anticipate that the structure of excited nuclei featuring an $\alpha$-cluster will be clearly manifested and further validated in future studies.
Additionally, findings from this work can serve as a reference for assessing whether a liquid-gas phase transition occurs within excited nucleus in $eA$ collisions. Should such a phase transition take place experimentally, an increased observation of light fragments is expected alongside minimal detection of heavy fragments. For very heavy excited nuclei, it is plausible that liquid-gas phase transition could occur in specific region where numerous light fragments are evaporated while other area undergoes fragmentation process or remaining smaller nucleus; this fragmentation process may deviate from traditional partitioning methods if significant $\alpha$-clustering is present.
Acknowledgments
The work of Shanxi Group was supported by the National Natural Science Foundation of China under Grant No. 12147215, the Shanxi Provincial Basic Research Program (Natural Science Foundation) under Grant No. 202103021224036, and the Fund for Shanxi “1331 Project” Key Subjects Construction. The work of K.K.O. was supported by the Agency of Innovative Development under the Ministry of Higher Education, Science and Innovations of the Republic of Uzbekistan within the fundamental project No. F3-20200929146 on analysis of open data on heavy-ion collisions at RHIC and LHC.
ORCID
Fu-Hu Liu https://orcid.org/$0$-0002-2261-6899
Khusniddin K. Olimov https://orcid.org/$0$-0002-1879-8458